# Constructing Safety Cases for AI Systems: A Reusable Template Framework

### v1

@@ Appendix B   List of Selected Studies
  • [S1] Simon Burton. A causal model of safety assurance for machine learning. preprint arXiv:2201.05451, 2022.
  • [S2] Daniella Tola and Peter Gorm Larsen. A co-simulation based approach for developing safety-critical systems. John Fitzgerald, Tomohiro Oda, and Hugo Daniel Macedo (Editors), 65, 2021.
@@
  • [S3] Alexander Rudolph, Stefan Voget, and Jürgen Mottok. A consistent safety case argumentation for artificial intelligence in safety related automotive systems. In ERTS 2018, 2018.
@@
  • [S4] Benjamin Herd, João-Vitor Zacchi, and Simon Burton. A deductive approach to safety assurance: Formalising safety contracts with subjective logic. In International Conference on Computer Safety, Reliability, and Security, pages 213–226. Springer, 2024.
@@
  • [S5] Yan Jia, Clare Verrill, Kieron White, Monica Dolton, Margaret Horton, Mufaddal Jafferji, and Ibrahim Habli. A deployment safety case for ai-assisted prostate cancer diagnosis. Computers in Biology and Medicine, 192:110237, 2025.
@@
  • [S6] Ewen Denney and Ganesh Pai. A formal basis for safety case patterns. In International Conference on Computer Safety, Reliability, and Security, pages 21–32. Springer, 2013.
@@
  • [S7] Jonas Krook, Yuvaraj Selvaraj, Wolfgang Ahrendt, and Martin Fabian. A formal-methods approach to provide evidence in automated-driving safety cases. preprint arXiv:2210.07798, 2022.
@@
  • [S8] Natsuki Hayama, Yoriyuki Yamagata, Hideaki Nishihara, and Yutaka Matsuno. A gsn-based requirement analysis of the eu AI regulation. In International Conference on Computer Safety, Reliability, and Security, pages 183–196. Springer, 2025.
@@
  • [S9] Peter Bishop and Robin Bloomfield. A methodology for safety case development. In Safety and Reliability, volume 20(1), pages 34–42. Taylor & Francis, 2000.
@@
  • [S10] Zoe Porter, Ibrahim Habli, John McDermid, and Marten Kaas. A principles-based ethics assurance argument pattern for AI and autonomous systems. AI and Ethics, 4(2):593–616, 2024.
@@
  • [S11] Oluwafemi Odu, Alvine Boaye Belle, Song Wang, and Kimya Khakzad Shahandashti. A prisma-driven bibliometric analysis of the scientific literature on assurance case patterns. preprint arXiv:2407.04961, 2024.
@@
  • [S12] Kimya Khakzad Shahandashti, Alvine B Belle, Timothy C Lethbridge, Oluwafemi Odu, and Mithila Sivakumar. A prisma-driven systematic mapping study on system assurance weakeners. Information and Software Technology, 175:107526, 2024.
@@
  • [S13] Magnus Gyllenhammar, Gabriel Rodrigues de Campos, and Martin Törngren. A safety argument fragment towards safe deployment of performant automated driving systems. In International Conference on Computer Safety, Reliability, and Security, pages 197–210. Springer, 2025.
@@
  • [S14] Ram Potham. A safety case for a deployed llm: Corrigibility as a singular target. In ILIAD 2: ODYSSEY, 2025.
@@
  • [S15] Ernest Wozniak, Carmen Cârlan, Esra Acar-Celik, and Henrik J Putzer. A safety case pattern for systems with machine learning components. In International Conference on Computer Safety, Reliability, and Security, pages 370–382. Springer, 2020.
@@
  • [S16] Tomek Korbak, Joshua Clymer, Benjamin Hilton, Buck Shlegeris, and Geoffrey Irving. A sketch of an AI control safety case. preprint arXiv:2501.17315, 2025.
@@
  • [S17] Ran Wei, Simon Foster, Haitao Mei, Fang Yan, Ruizhe Yang, Ibrahim Habli, Colin O’Halloran, Nick Tudor, Tim Kelly, and Yakoub Nemouchi. Access: Assurance case centric engineering of safety–critical systems. Journal of Systems and Software, 213:112034, 2024.
@@
  • [S18] Simon Burton and Benjamin Herd. Addressing uncertainty in the safety assurance of machine-learning. Frontiers in Computer Science, 5:1132580, 2023.
@@
  • [S19] Victoria J Hodge and Matt Osborne. Agile development for safety assurance of machine learning in autonomous systems (agileamlas). Array, page 100482, 2025.
@@
  • [S20] Bernhard Kaiser, Michael Soden, Rüdiger Diefenbach, and Eckhardt Holz. An agile approach to safety cases for autonomous systems through model-based engineering and simulation. In Safety-Critical Systems, pages 1–23, 02 2025.
@@
  • [S21] Marie Davidsen Buhl, Jacob Pfau, Benjamin Hilton, and Geoffrey Irving. An alignment safety case sketch based on debate. preprint arXiv:2505.03989, 2025.
@@
  • [S22] Francis Rhys Ward and Ibrahim Habli. An assurance case pattern for the interpretability of machine learning in safety-critical systems. In International Conference on Computer Safety, Reliability, and Security, pages 395–407. Springer, 2020.
@@
  • [S23] Joshua Clymer, Jonah Weinbaum, Robert Kirk, Kimberly Mai, Selena Zhang, and Xander Davies. An example safety case for safeguards against misuse. preprint arXiv:2505.18003, 2025.
@@
  • [S24] Sunil Nair, Jose Luis De La Vara, Mehrdad Sabetzadeh, and Lionel Briand. An extended systematic literature review on provision of evidence for safety certification. Information and Software Technology, 56(7):689–717, 2014.
@@
  • [S25] Uma D Ferrell and Alfred H Andy Anderegg. Applicability of ul 4600 to unmanned aircraft systems (uas) and urban air mobility (uam). In 2020 AIAA/IEEE 39th Digital Avionics Systems Conference (DASC), pages 1–7. IEEE, 2020.
@@
  • [S26] João Paulo Costa de Araujo, Balahari Vignesh Balu, Eik Reichmann, Jessica Kelly, Stefan Kugele, Núria Mata, and Lars Grunske. Applying concept-based models for enhanced safety argumentation. In 2024 IEEE 35th International Symposium on Software Reliability Engineering (ISSRE), pages 272–283. IEEE, 2024.
@@
  • [S27] Tim P Kelly. Arguing safety-a systematic approach to safety case management. DPhil Thesis York University, Department of Computer Science Report YCST, 1999.
@@
  • [S28] Stephen Barrett, Philip Fox, Joshua Krook, Tuneer Mondal, Simon Mylius, and Alejandro Tlaie. Assessing confidence in frontier AI safety cases. preprint arXiv:2502.05791, 2025.
@@
  • [S29] Ramneet Kaur, Radoslav Ivanov, Matthew Cleaveland, Oleg Sokolsky, and Insup Lee. Assurance case patterns for cyber-physical systems with deep neural networks. In International Conference on Computer Safety, Reliability, and Security, pages 82–97. Springer, 2020.
@@
  • [S30] Thomas Chowdhury. Assurance Case Templates: Principles for Their Development and Criteria for Their Evaluation. PhD thesis, McMaster University, Hamilton, Ontario, Canada, 2021.
@@
  • [S31] Vincent Mussot, Eric Jenn, Florent Chenevier, Ramon Conejo Laguna, Yassir Id Messaoud, Jean-Loup Farges, Anthony Fernandes Pires, Florent Latombe, and Stephen Creff. Assurance cases to face the complexity of ml-based systems verification. In Embedded Real Time System Congress, ERTS’24, 2024.
@@
  • [S32] Robin Bloomfield and John Rushby. Assurance of AI systems from a dependability perspective. preprint arXiv:2407.13948, 2024.
@@
  • [S33] Ewen Denney and Ganesh Pai. Assurance-driven design of machine learning-based functionality in an aviation systems context. In IEEE/AIAA 42nd Digital Avionics Systems Conference (DASC), pages 1–10. IEEE, 2023.
@@
  • [S34] Andrzej Wardziński and Aleksander Jarzębowicz. Automated generation of modular assurance cases with the system assurance reference model. Formal Aspects of Computing, 36(4):1–29, 2024.
@@
  • [S35] Kimya Khakzad Shahandashti, Song Wang, Oluwafemi Odu, and Alvine Boaye Belle. Automatic instantiation of assurance cases from patterns using large language models. In IEEE/ACM 4th International Conference on AI Engineering –Software Engineering for AI (CAIN), pages 156–167. IEEE, 2025.
  • [S36] Julian Münch, Michael Vierhauser, and Timo Kehrer. Automating safety argument change impact analysis for machine learning components. Empirical Software Engineering, 29(1):1–31, 2024.
  • [S37] Joshua Smith, David Krueger, and Zoe Porter. Balancing the risks and benefits of using large language models to support assurance case development. preprint arXiv:2505.11312, 2025.
@@
  • [S38] Torben Nass. Checkable Safety Arguments: A Modeling Framework Supporting the Maintenance of Safety Arguments Consistent with System Development Artifacts. PhD thesis, University of York, 2022.
  • [S39] Richard Hawkins and Ibrahim Habli. Confidence assessment in safety argument structure: Quantitative vs qualitative approaches. Reliability Engineering & System Safety, 167:479–493, 2017.
@@
  • [S40] Esra Acar-Celik. Design and automatic generation of safety cases of ML-enabled autonomous driving systems. PhD thesis, Technical University of Munich, 2023.
@@
  • [S41] Jinjun Shan and Alvine Boaye Belle. Design of the safety case of the reinforcement learning-enabled component of a quanser autonomous vehicle. preprint arXiv:2405.02021, 2024.
  • [S42] Julian Flad, Stefan Kugele, Markus Reischl, and Alois Knoll. Design-time specification of dynamic modular safety cases in support of run-time safety assessment. In International Conference on Computer Safety, Reliability, and Security, pages 183–195. Springer, 2022.
  • [S43] Peter Bishop and Robin Bloomfield. Developing compelling safety cases. Safecomp 1998: Computer Safety, Reliability and Security, pages 21–36, 1998.
  • [S44] Ibrahim Habli, Tim Kelly, and Mark A Sujan. Dynamic assurance cases: A pathway to trusted autonomy. Safety Science, 128:104745, 2020.
  • [S45] Richard Hawkins, Ibrahim Habli, and Tim Lawton. Dynamic safety assurance of autonomous cyber-physical systems. preprint arXiv:2403.05641, 2024.
  • [S46] David Krueger, Zoe Porter, and Ibrahim Habli. Dynamic safety cases for frontier ai. preprint arXiv:2504.07115, 2025.
  • [S47] Ernest Wozniak, Richard Heinrich, and Julian Reich. Ergo, smirk is safe: A safety case for a machine learning component in a pedestrian automatic emergency brake system. preprint arXiv:2312.00690, 2023.
@@
  • [S48] Alan FT Winfield, Marina Jirotka, Max Gogoshin, Gonenc Gürkaynak, and Vivek Nallur. Ethical assurance: A practical approach to the responsible design, development, and deployment of data-driven technologies. AI and Ethics, 2(3):349–365, 2022.
  • [S49] Alan Burns, Iain Bate, and Tim Kelly. Evaluating ai-based components in autonomous railway systems. IEEE Transactions on Intelligent Transportation Systems, 2023.
@@
  • [S50] Oluwafemi Odu, Song Wang, Alvine Boaye Belle, and Kimya Khakzad Shahandashti. Explainable compliance detection with multi-hop natural language inference on assurance case structure. preprint arXiv:2409.05711, 2024.
@@
  • [S51] Mithila Sivakumar, Alvine Boaye Belle, Jinjun Shan, and Kimya Khakzad Shahandashti. Exploring the capabilities of large language models for the generation of safety cases: The case of gpt-4. preprint arXiv:2405.02493, 2024.
  • [S52] Ewen Denney and Ganesh Pai. Generation of safety case argument-fragments from safety contracts. In NASA Formal Methods Symposium, pages 64–78. Springer, 2014.
@@
  • [S53] Mithila Sivakumar, Alvine Boaye Belle, Jinjun Shan, and Kimya Khakzad Shahandashti. Gpt-4 and safety case generation: An exploratory analysis. preprint arXiv:2406.01976, 2024.
@@
  • [S54] Abhishek Goswami, Subhrajit Maiti, Nilotpal Sinha, and Himadri Sekhar Das. Graphical safety assurance case using goal structuring notation (gsn): Challenges, opportunities and a framework for autonomous trains. In International Conference on Computer Safety, Reliability, and Security, pages 168–181. Springer, 2022.
@@
  • [S55] Ewen Denney and Ganesh Pai. Guided integration of formal verification in assurance cases. Innovations in Systems and Software Engineering, 10(4):323–334, 2014.
@@
  • [S56] Jamal Mouhawesh, Irfan Sljivo, Barbara Gallina, and Ibrahim Habli. Identifying run-time monitoring requirements for autonomous systems through the analysis of safety arguments. In International Conference on Computer Safety, Reliability, and Security, pages 70–82. Springer, 2019.
  • [S57] Harold Thimbleby, Patrick Oladimeji, and Paul Curzon. Introducing safety cases for health it. Journal of the American Medical Informatics Association, 27(4):621–627, 2020.
  • [S58] Tim Kelly and John McDermid. Issues in the application of software safety standards. Safety Science, 24(2):77–89, 1996.
  • [S59] Daniel Hirsch, Joshua Clymer, Ibrahim Habli, Zoe Porter, and Marten Kaas. Justified evidence collection for argument-based AI fairness assurance. preprint arXiv:2504.00441, 2025.
  • [S60] Carlos Rodriguez and Tim Kelly. Leveraging artifact trees to evolve and reuse safety cases. In International Conference on Computer Safety, Reliability, and Security, pages 157–170. Springer, 2021.
  • [S61] Oluwafemi Odu, Alvine Boaye Belle, Kimya Khakzad Shahandashti, and Song Wang. Llm-based automated mitigation and assurance case generation against threats to AI systems. preprint arXiv:2505.09807, 2025.
  • [S62] Zhiwei Huang, Yuan Zhao, Wei Xu, Meng Wang, and Lichao Sun. Llm-based safety case generation for baidu apollo: Are we there yet? preprint arXiv:2504.03989, 2025.
@@
  • [S63] Simon Burton, Lydia Gauerhof, and Christian Heinzemann. Making the case for safety of machine learning in highly automated driving. SAE International Journal of Connected and Automated Vehicles, 2(1):219–230, 2019.
@@
  • [S64] Alex Whitworth and Jonathan Borrie. Manipulation attacks by misaligned ai: Risk analysis and safety case framework. preprint arXiv:2506.00310, 2025.
  • [S65] Bassel Rajab, Fernand Belmonte, Radu Calinescu, and Ganesh Pai. Measure confidence of assurance cases in safety-critical domains. preprint arXiv:2209.05506, 2022.
  • [S66] David Pumfrey, Ibrahim Habli, and John McDermid. Measuring safety: applying psm to the system safety domain. In Adelard Workshop on Assurance Cases, 2003.
  • [S67] Alan Burns, Patrick Graydon, and Lorenzo Strigini. Modalas: addressing assurance for learning-enabled autonomous systems in the face of uncertainty. In International Conference on Computer Safety, Reliability, and Security, pages 241–252. Springer, 2024.

### v3

@@ Appendix B   List of Selected Studies
  • [S1] S. Burton, “A causal model of safety assurance for machine learning,” arXiv preprint arXiv:2201.05451, 2022.
  • [S2] D. Tola and P. G. Larsen, “A co-simulation based approach for developing safety-critical systems,” in Proceedings of the 18th International Overture Workshop, vol. 65, 2021.
@@
  • [S3] A. Rudolph, S. Voget, and J. Mottok, “A consistent safety case argumentation for artificial intelligence in safety related automotive systems,” in ERTS 2018, 2018.
@@
  • [S4] B. Herd, J.-V. Zacchi, and S. Burton, “A deductive approach to safety assurance: Formalising safety contracts with subjective logic,” in International Conference on Computer Safety, Reliability, and Security. Springer Nature Switzerland Cham, 2024, pp. 213–226.
@@
  • [S5] Y. Jia, C. Verrill, K. White, M. Dolton, M. Horton, M. Jafferji, and I. Habli, “A deployment safety case for ai-assisted prostate cancer diagnosis,” Computers in biology and medicine, vol. 192, p. 110237, 2025.
@@
  • [S6] E. Denney and G. Pai, “A formal basis for safety case patterns,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2013, pp. 21–32.
@@
  • [S7] J. Krook, Y. Selvaraj, W. Ahrendt, and M. Fabian, “A formal-methods approach to provide evidence in automated-driving safety cases,” arXiv preprint arXiv:2210.07798, 2022.
@@
  • [S8] N. Hayama, Y. Yamagata, H. Nishihara, and Y. Matsuno, “A gsn-based requirement analysis of the eu AI regulation,” in International Conference on Computer Safety, Reliability, and Security. Springer Nature Switzerland Cham, 2025, pp. 183–196.
@@
  • [S9] P. Bishop and R. Bloomfield, “A methodology for safety case development,” in Safety and Reliability, vol. 20, no. 1. Taylor & Francis, 2000, pp. 34–42.
@@
  • [S10] Z. Porter, I. Habli, J. McDermid, and M. Kaas, “A principles-based ethics assurance argument pattern for AI and autonomous systems,” AI and Ethics, vol. 4, no. 2, pp. 593–616, 2024.
@@
  • [S11] O. Odu, A. B. Belle, S. Wang, and K. K. Shahandashti, “A prisma-driven bibliometric analysis of the scientific literature on assurance case patterns,” arXiv preprint arXiv:2407.04961, 2024.
@@
  • [S12] K. K. Shahandashti, A. B. Belle, T. C. Lethbridge, O. Odu, and M. Sivakumar, “A prisma-driven systematic mapping study on system assurance weakeners,” Information and Software Technology, vol. 175, p. 107526, 2024.
@@
  • [S13] M. Gyllenhammar, G. R. d. Campos, and M. Törngren, “A safety argument fragment towards safe deployment of performant automated driving systems,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2025, pp. 197–210.
@@
  • [S14] R. Potham, “A safety case for a deployed llm: Corrigibility as a singular target via debate,” 2025.
@@
  • [S15] E. Wozniak, C. Cârlan, E. Acar-Celik, and H. J. Putzer, “A safety case pattern for systems with machine learning components,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2020, pp. 370–382.
@@
  • [S16] T. Korbak, J. Clymer, B. Hilton, B. Shlegeris, and G. Irving, “A sketch of an AI control safety case,” arXiv preprint arXiv:2501.17315, 2025.
@@
  • [S17] R. Wei, S. Foster, H. Mei, F. Yan, R. Yang, I. Habli, C. O’Halloran, N. Tudor, T. Kelly, and Y. Nemouchi, “Access: Assurance case centric engineering of safety–critical systems,” Journal of Systems and Software, vol. 213, p. 112034, 2024.
@@
  • [S18] S. Burton and B. Herd, “Addressing uncertainty in the safety assurance of machine-learning,” Frontiers in computer science, vol. 5, p. 1132580, 2023.
@@
  • [S19] V. J. Hodge and M. Osborne, “Agile development for safety assurance of machine learning in autonomous systems (agileamlas),” Array, p. 100482, 2025.
@@
  • [S20] B. Kaiser, M. Soden, R. Diefenbach, and E. Holz, “An agile approach to safety cases for autonomous systems through model-based engineering and simulation,” in Proc. 33rd Saf. Crit. Syst. Symp, 2025.
@@
  • [S21] M. D. Buhl, J. Pfau, B. Hilton, and G. Irving, “An alignment safety case sketch based on debate,” arXiv preprint arXiv:2505.03989, 2025.
@@
  • [S22] F. R. Ward and I. Habli, “An assurance case pattern for the interpretability of machine learning in safety-critical systems,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2020, pp. 395–407.
@@
  • [S23] J. Clymer, J. Weinbaum, R. Kirk, K. Mai, S. Zhang, and X. Davies, “An example safety case for safeguards against misuse,” arXiv preprint arXiv:2505.18003, 2025.
@@
  • [S24] S. Nair, J. L. De La Vara, M. Sabetzadeh, and L. Briand, “An extended systematic literature review on provision of evidence for safety certification,” Information and Software Technology, vol. 56, no. 7, pp. 689–717, 2014.
@@
  • [S25] U. D. Ferrell and A. H. A. Anderegg, “Applicability of ul 4600 to unmanned aircraft systems (uas) and urban air mobility (uam),” in 2020 AIAA/IEEE 39th Digital Avionics Systems Conference (DASC). IEEE, 2020, pp. 1–7.
@@
  • [S26] J. P. C. de Araujo, B. V. Balu, E. Reichmann, J. Kelly, S. Kugele, N. Mata, and L. Grunske, “Applying concept-based models for enhanced safety argumentation,” in 2024 IEEE 35th International Symposium on Software Reliability Engineering (ISSRE). IEEE, 2024, pp. 272–283.
@@
  • [S27] T. P. Kelly, “Arguing safety-a systematic approach to safety case management,” Department of Computer Science, The University of York, 1998.
@@
  • [S28] S. Barrett, P. Fox, J. Krook, T. Mondal, S. Mylius, and A. Tlaie, “Assessing confidence in frontier AI safety cases,” arXiv preprint arXiv:2502.05791, 2025.
@@
  • [S29] R. Kaur, R. Ivanov, M. Cleaveland, O. Sokolsky, and I. Lee, “Assurance case patterns for cyber-physical systems with deep neural networks,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2020, pp. 82–97.
@@
  • [S30] T. Chowdhury, “Assurance case templates: Principles for their development and criteria for their evaluation,” Ph.D. dissertation, McMaster University, 2021.
@@
  • [S31] V. Mussot, E. Jenn, F. Chenevier, R. C. Laguna, Y. I. Messaoud, J.-L. Farges, A. F. Pires, F. Latombe, and S. Creff, “Assurance cases to face the complexity of ml-based systems verification,” in Embedded Real Time System Congress, ERTS’24, 2024.
@@
  • [S32] R. Bloomfield and J. Rushby, “Assurance of AI systems from a dependability perspective,” arXiv preprint arXiv:2407.13948, 2024.
@@
  • [S33] E. Denney and G. Pai, “Assurance-driven design of machine learning-based functionality in an aviation systems context,” in 2023 IEEE/AIAA 42nd Digital Avionics Systems Conference (DASC). IEEE, 2023, pp. 1–10.
@@
  • [S34] A. Wardziński and A. Jarzębowicz, “Automated generation of modular assurance cases with the system assurance reference model,” Formal Aspects of Computing, vol. 36, no. 4, pp. 1–29, 2024.
@@
  • [S35] O. Odu, A. B. Belle, S. Wang, S. Kpodjedo, T. C. Lethbridge, and H. Hemmati, “Automatic instantiation of assurance cases from patterns using large language models,” Journal of Systems and Software, vol. 222, p. 112353, 2025.
  • [S36] C. Cârlan, L. Gauerhof, B. Gallina, and S. Burton, “Automating safety argument change impact analysis for machine learning components,” in 2022 IEEE 27th Pacific Rim International Symposium on Dependable Computing (PRDC). IEEE, 2022, pp. 43–53.
  • [S37] S. Diemert, E. Cyffka, N. Anwari, O. Foster, T. Viger, L. Millet, and J. Joyce, “Balancing the risks and benefits of using large language models to support assurance case development,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2025, pp. 209–225.
@@
  • [S38] C. Cârlan, “Checkable safety arguments-a modeling framework supporting the maintenance of safety arguments consistent with system development artifacts,” Ph.D. dissertation, Technische Universität München, 2025.
  • [S39] Y. Idmessaoud, D. Dubois, and J. Guiochet, “Confidence assessment in safety argument structure-quantitative vs. qualitative approaches,” International Journal of Approximate Reasoning, vol. 165, p. 109100, 2024.
@@
  • [S40] M. Sivakumar, “Design and automatic generation of safety cases of ml-enabled autonomous driving systems,” Master’s thesis, York University, 2024.
@@
  • [S41] M. Sivakumar, A. B. Belle, J. Shan, O. Odu, and M. Yuan, “Design of the safety case of the reinforcement learning-enabled component of a quanser autonomous vehicle,” in 2024 IEEE 32nd International Requirements Engineering Conference Workshops (REW). IEEE, 2024, pp. 57–67.
  • [S42] E. Mirzaei, C. CARLAN, C. Thomas, and B. Gallina, “Design-time specification of dynamic modular safety cases in support of run-time safety assessment,” in Proceedings of the 30th Safety-Critical Systems Symposium (SCSC), 2022.
  • [S43] R. Hawkins, “Developing compelling safety cases,” arXiv preprint arXiv:2502.00911, 2025.
  • [S44] E. Asaadi, E. Denney, J. Menzies, G. J. Pai, and D. Petroff, “Dynamic assurance cases: a pathway to trusted autonomy,” Computer, vol. 53, no. 12, pp. 35–46, 2020.
  • [S45] S. Ramakrishna, “Dynamic safety assurance of autonomous cyber-physical systems,” Ph.D. dissertation, Vanderbilt University, 2022.
  • [S46] C. Cârlan, F. Gomez, Y. Mathew, K. Krishna, R. King, P. Gebauer, and B. R. Smith, “Dynamic safety cases for frontier ai,” arXiv preprint arXiv:2412.17618, 2024.
  • [S47] M. Borg, J. Henriksson, K. Socha, O. Lennartsson, E. Sonnsjö Lönegren, T. Bui, P. Tomaszewski, S. R. Sathyamoorthy, S. Brink, and M. Helali Moghadam, “Ergo, smirk is safe: a safety case for a machine learning component in a pedestrian automatic emergency brake system,” Software quality journal, vol. 31, no. 2, pp. 335–403, 2023.
@@
  • [S48] C. Burr and D. Leslie, “Ethical assurance: a practical approach to the responsible design, development, and deployment of data-driven technologies,” AI and Ethics, vol. 3, no. 1, pp. 73–98, 2023.
  • [S49] J. Roßbach, O. De Candido, A. Hammam, and M. Leuschel, “Evaluating ai-based components in autonomous railway systems: A methodology,” in German Conference on Artificial Intelligence (Künstliche Intelligenz). Springer, 2024, pp. 190–203.
@@
  • [S50] F. Ikhwantri and D. Marijan, “Explainable compliance detection with multi-hop natural language inference on assurance case structure,” arXiv preprint arXiv:2506.08713, 2025.
@@
  • [S51] M. Sivakumar, A. B. Belle, J. Shan, and K. K. Shahandashti, “Exploring the capabilities of large language models for the generation of safety cases: the case of gpt-4,” in 2024 IEEE 32nd International Requirements Engineering Conference Workshops (REW). IEEE, 2024, pp. 35–45.
  • [S52] I. Sljivo, B. Gallina, J. Carlson, and H. Hansson, “Generation of safety case argument-fragments from safety contracts,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2014, pp. 170–185.
@@
  • [S53] M. Sivakumar, A. B. Belle, J. Shan, and K. K. Shahandashti, “Gpt-4 and safety case generation: An exploratory analysis,” arXiv preprint arXiv:2312.05696, 2023.
@@
  • [S54] M. Chelouati, A. Boussif, J. Beugin, and E.-M. El Koursi, “Graphical safety assurance case using goal structuring notation (gsn)—challenges, opportunities and a framework for autonomous trains,” Reliability Engineering & System Safety, vol. 230, p. 108933, 2023.
@@
  • [S55] I. Sljivo, E. Denney, and J. Menzies, “Guided integration of formal verification in assurance cases,” in International Conference on Formal Engineering Methods. Springer, 2023, pp. 172–190.
@@
  • [S56] R. Hawkins and P. Ryan Conmy, “Identifying run-time monitoring requirements for autonomous systems through the analysis of safety arguments,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2023, pp. 11–24.
  • [S57] G. Despotou, S. White, T. Kelly, and M. Ryan, “Introducing safety cases for health it,” in 2012 4th International Workshop on Software Engineering in Health Care (SEHC). IEEE, 2012, pp. 44–50.
  • [S58] M. J. Squair, “Issues in the application of software safety standards,” in ACM International Conference Proceeding Series, vol. 162, 2006, pp. 13–26.
  • [S59] A. Sabuncuoglu, C. Burr, and C. Maple, “Justified evidence collection for argument-based AI fairness assurance,” in Proceedings of the 2025 ACM Conference on Fairness, Accountability, and Transparency, 2025, pp. 18–28.
  • [S60] A. Agrawal, S. Khoshmanesh, M. Vierhauser, M. Rahimi, J. Cleland-Huang, and R. Lutz, “Leveraging artifact trees to evolve and reuse safety cases,” in 2019 IEEE/ACM 41st International Conference on Software Engineering (ICSE). IEEE, 2019, pp. 1222–1233.
  • [S61] Y. Fujiwara, T. Tuchida, R. Miyata, H. Washizaki, and N. Ubayashi, “Llm-based automated mitigation and assurance case generation against threats to AI systems,” in 2025 IEEE Conference on Artificial Intelligence (CAI). IEEE, 2025, pp. 906–911.
  • [S62] O. Odu, A. B. Belle, and S. Wang, “Llm-based safety case generation for baidu apollo: Are we there yet?” in 2025 IEEE/ACM 4th International Conference on AI Engineering–Software Engineering for AI (CAIN). IEEE, 2025, pp. 222–233.
@@
  • [S63] S. Burton, L. Gauerhof, and C. Heinzemann, “Making the case for safety of machine learning in highly automated driving,” in International Conference on Computer Safety, Reliability, and Security. Springer, 2017, pp. 5–16.
@@
  • [S64] R. Dassanayake, M. Demetroudi, J. Walpole, L. Lentati, J. R. Brown, and E. J. Young, “Manipulation attacks by misaligned ai: Risk analysis and safety case framework,” arXiv preprint arXiv:2507.12872, 2025.
  • [S65] C.-L. Lin, W. Shen, S. Drager, and B. Cheng, “Measure confidence of assurance cases in safety-critical domains,” in Proceedings of the 40th International Conference on Software Engineering: New Ideas and Emerging Results, 2018, pp. 13–16.
  • [S66] J. Murdoch, G. Clark, A. Powell, and P. Caseley, “Measuring safety: applying psm to the system safety domain,” in Proceedings of the 8th Australian workshop on Safety critical systems and software-Volume 33, 2003, pp. 47–55.
  • [S67] M. A. Langford, K. H. Chan, J. E. Fleck, P. K. McKinley, and B. H. Cheng, “MoDALAS: addressing assurance for learning-enabled autonomous systems in the face of uncertainty,” Software and systems modeling, vol. 22, no. 5, pp. 1543–1563, 2023.